Consider the following game: - 2 LR 2

a.) Find all pure-strategy Nash equilibria.

b.) *Find all mixed-strategy Nash equilibria.

c.) Explain why, in any mixed-strategy equilibrium, each player must be indifferent between the pure strategies that she randomizes over.
Answer #1

(1-9) l pulsu (1) D4?s ? (as Pure Strategy Nash equilibrium No fure strategy hash equilibrium. (b) der Player 1 Choose loo sa29 = 5.4 = q = 1 2 (1-9) = 1/2 Foo Player-2 expected Payoff from a = expected Payoff from R 2p+ 3(1-P) = 3b + 2(1-6) 26+ 3-3p(e, consider hayer-1. exlected Payoff from u = 9+4 11 D= 5-9 so if Elul > E(0) (that is he is not indifJerent ) heb 9ч » S-4